Q8. Given the following data: Lead Time Beginning Lot Size LAL LAL LAL Usage per (weeks) 0 30 70 The end product A is made from components B and C. The authorized MPS calls for 30 A to be n ? completed in week 5. What will be the Net requirement for Item A. 70 B. 120 C. 80 D. 30 E. 150 Q9. Th ost per unit is $10.09 at the best operating level. e unit cost will when the output is less than the output at best operating level, if the A. >$10.09 B. <$10.09 C. = $10.09 D. Both A&B E. Both B&C One of the basic purposes of inventory analysis in manufacturing and stockkeeping services is to rmine the level of quality to specify. False YZ medical facility provides medical services to patients in Russelville. The design capacity atients per hour, and the effective capacity is 35 patients per hour. Yesterday the medical worked for 7 hours and served 200 patients. What is the effective utilization? 5.23%

Can anyone please help me solving these problems ASAP?
Answer #1

Q8. C. 80

Net requirement of C = Net requirement of Parent item (A) * Usage per parent - Beginning inventory of C

= 30*5 - 70

= 80.

Q9. A. >$10.09

When output is less, then labor cost and fixed cost per unit increases

Q10. False

Basic purpose of inventory analysis is to determine optimal inventory level

Q11. 81.63 %

Effective utilization = Actual output / Effective capacity = 200 / (35*7) = 81.63 %